2 11 4. Let C = Show that C2 – 4C +31 = 0. Use this equation to calculate C-1 12
Answer #1

< (?). (?!).(??) Now, - 96+33 + (115)-(41)+(33) 4 & > ise. ^-4631 = 0 => c (C-41) = -32 => -14 (6-98)= 6? [ multiplying e bo